OpenStudy (anonymous):

proooooof the\[\beta ^{th}\] projection map \[\pi _{\beta} : \Pi _{\alpha \in \Delta } \rightarrow X _{\beta}\] is continuous and open but not be closed

OpenStudy (perl):

hmm, is there a definition for projection map? i only know projection of a function , like projection f(x,y)= (0,y)

OpenStudy (anonymous):

nooooo

OpenStudy (anonymous):

projection map not like projection of a function

OpenStudy (perl):

sorry i dont know what that is. what class is this? or book

OpenStudy (anonymous):

this is the Luther of the book paul e. long

OpenStudy (anonymous):

oooooh here its like \[\Pi _{\alpha \in \Delta } X _{\alpha} \rightarrow X _{\beta}\]

OpenStudy (anonymous):

Ah, got it

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

So, if U is open in X_beta, then \[\pi_\beta^{-1}(U) = U \times \Pi_{\alpha \in \Delta\backslash\beta}\], which is in the basis according to the definition of a product topology.

OpenStudy (anonymous):

So, that satisfies the definition for continuity. (That Latex was so hard!)

OpenStudy (anonymous):

how to proof if it open

OpenStudy (anonymous):

what about need not to be closed

OpenStudy (anonymous):

Oh, whoops. Let me work on those

OpenStudy (anonymous):

where the continuous proof and you know that projection map is onto but not 1-1

OpenStudy (anonymous):

Ok, so, just take an arbitrary element in the basis of the product topology (so, a union of inverse projections of open sets from the X's). Then apply some projection map to it. If the projection map goes to one of the X's that contributed an inverse projection to that basis element, then the map goes back to the original open set, and it's open. Otherwise, the map goes back to an entire X, which is also open. Either way, the projection maps basis elements to open sets, so the entire map is open.

OpenStudy (anonymous):

It need not be closed because you could go through the same argument for a basis of closed sets, but then when you map back to an entire X, that's an open set, so the argument doesn't hold.

OpenStudy (anonymous):

but what will happens if closure \[ of \left( X _{\beta} \right)\]

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

πβ:Πα∈Δ→ closure (Xβ)

OpenStudy (anonymous):

I'm not sure what you mean...you can only take the closure of subset, not the entire topology.

OpenStudy (anonymous):

but it is in product spaces

OpenStudy (anonymous):

Oh, I guess I see what you mean. But the projection still maps to the set itself, not its closure.

OpenStudy (anonymous):

but what if it be in closure dose the the projection map will find some elements to map

OpenStudy (anonymous):

Sorry, not sure what you're saying?